Find a root of the equation 3x^3+10x^2+10x+7 = 0 which is close to −2.0, using
the Berge -Vieta method. Perform two iterations of the method.
1
Expert's answer
2020-04-28T14:48:14-0400

At the first step of any iteration of the Birge-Vieta method, we write down the coefficients of the equation and put that

b0=a0;

c0=a0.

Also, we have the initial approximation p0=-2.

Second step: find a coefficient


"b_i=a_i+p_0\u22c5b_{i-1}."

Third step: find


"c_i=b_i+p_0\u22c5c_{i-1}."

Fourth step: find


"p_1=p_0-\frac{b_n}{c_{i-1}}."

Fifth step: when


"|p_1-p_0|≤\epsilon,"


we found the answer.


It is better to arrange the coefficients in the form of a table. Therefore, at the first iteration we have



and


"p_1=p_0-\frac{b_3}{c_2}=-2-\frac{3}{6}=-2.5."

The second iteration gives



and


"p_2=p_1-\frac{b_3}{c_2}=-2.5-\frac{-2.375}{16.25}=-2.3539."